Let f : R²→R,f(x,y) =2xy/x²+y² When (x,y) ≠(0,0) and f(x,y) =0, otherwise. How to Check whether or not f has a directional derivative at (0, 0) in the direction θ=π/4? Deduce that the function f is not differentiable at the point (0, 0)?

1
Expert's answer
2021-07-27T09:02:51-0400

f : R²→R,f(x, y) = "\dfrac{2xy}{x^2 + y^2}"


We will first find out "\nabla" ( f )


"\nabla" ( f ) = "\dfrac{2y^3 - 2x^2y}{(x^2 + y^2)^2} i + \dfrac{2x^3 - 2xy^2}{(x^2 + y^2)^2} j"



Directional derivative is given by "\nabla" ( f ) . "\overrightarrow{u}"


Where "\overrightarrow{u}" = "cos \dfrac{\pi}{4} i + sin\dfrac{\pi}{4} j" = "\dfrac{1}{\sqrt{2}} i + \dfrac{1}{\sqrt{2}} j"



"\nabla" ( f ) . "\overrightarrow{u}" = "\dfrac{\sqrt{2}y^3 - \sqrt{2}x^2y}{(x^2 + y^2)^2} + \dfrac{\sqrt{2}x^3 - \sqrt{2}xy^2}{(x^2 + y^2)^2}"


Now as x "\rightarrow" 0 and y "\rightarrow" 0 "\nabla" ( f ) . "\overrightarrow{u}" does not exist. Hence, the directional derivative of f does not exist at x = 0 and y = 0.



f'(x, y) = "\dfrac{∂f}{∂x} + \dfrac{∂f}{∂y}" = "\dfrac{2y^3 - 2x^2y}{(x^2 + y^2)^2} + \dfrac{2x^3 - 2xy^2}{(x^2 + y^2)^2}"


Now as x "\rightarrow" 0 and y "\rightarrow" 0 f'(x, y) does not exist. Hence, the derivative of f does not exist at x = 0 and y = 0.